Operations Report – December 4th

SOL: 7

Name of person filing report: Ashley Kowalski

Non-nominal systems: N/A

Notes on non-nominal systems: N/A


Spirit rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: ~270 gallons remaining
Water (static tank): The crew was informed that one load of water resupply was delivered today (120 gallons). Yesterday, we were told we had 175 gallons left. Today, we estimate we used ~25 gallons. So, before the water resupply, we likely had ~150 gallons. Thus, once the resupply was delivered, we now have about ~270 gallons.
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The Crew Engineer and Health and Safety Officer performed the suggestions provided to the crew by Mission Support. We realized that the problem actually was with the power strip that all of the chargers were plugged into. This is the process of elimination that we went through in order to get to that conclusion:

We first opened EVA suits #3, #4, and #6 and inspected the wiring and the fuses. All of the wiring looked good. The Crew Engineer replaced the blown fuse in EVA Suit #4. We also confirmed that all of the chargers worked. They all read >13 V.

Next, we swapped batteries between suits. We took a battery from Suit #1, one of the known functional suits, and put it into each of the non-functioning suits (#3, #4, and #6). When we did this, each one of the suits turned on and the air flowed normally. Thus, we could conclude that there was nothing wrong with the suits themselves, but rather the problem lay with the batteries we took out of those suits.

Then, we put each of the batteries from the non-functioning suits into Suit #1 (the functioning suit) to see if those batteries would increase from their low values to a value >12 V. The interesting part here is that the battery would not increase its charge even in the good suit. In fact, it was losing charge. So, we still thought the problem must be with the battery.

However, we decided to try one last thing. We thought that perhaps the problem was with the power strip that all the chargers were plugged into. The lights on those chargers did not consistently shine red or green (depending on if the battery was charger or fully charged), so we plugged in EVA suits #3, #4, and #6 with their original batteries (the ones that were reading low voltage values) into a different power strip that was found on the table shelf directly under where the two-piece suits stand and left them to charge for a bit. When we came back, the battery for suit #3 read 13.3 V, the battery for suit #4 read 13.76 V, and the battery for suit #6 read 12.33 V.

Thus, we can confidently conclude that the power strip was the actual issue. All two-piece suits now have charged, are reading the correct voltage values, and are ready for use. Please let us know what you would like us to do with the power strip having issues. We have labeled it indicating that it is having issues.

Finally, the two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Additional information can be found in the GreenHab Report.
WATER USE: ~34 gallons (~3 gallons today)
Heater: On
Supplemental light: Lights were checked at 11:15 AM, and appear OFF as expected. The GreenHab Officer will check that they are on after 10 PM tonight.
Harvest: N/A

Summary of ScienceDome operations: No operations performed today by the crew in the Science Dome.
Dual split: N/A

Summary of RAM operations: The multimeter from the RAM was used to check the voltages of the two-piece EVA suits during today’s troubleshooting/inspections.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: The odor from the toilet area has dissipated. There is no longer an issue with smell.

Copyright © The Mars Society. All rights reserved. | Main Site